# MagicDriveDiT: High-Resolution Long Video Generation for Autonomous Driving with Adaptive Control

### v2

@@ Abstract
The rapid advancement of diffusion models has greatly improved video synthesis, especially in controllable video generation, which is vital for applications like autonomous driving. Although DiT with 3D VAE has become a standard framework for video generation, it introduces challenges in controllable driving video generation, especially for geometry control, rendering existing control methods ineffective. To address these issues, we propose MagicDriveDiT, a novel approach that integrates the MVDiT block and spatial-temporal conditional encoding to enable multi-view video generation and precise geometric control. Additionally, we introduce an efficient method for obtaining contextual descriptions for videos to support diverse textual control, along with a progressive training strategy using mixed video data to enhance training efficiency and generalizability. Consequently, MagicDriveDiT enables multi-view driving video synthesis with $\times3$ resolution and 4×4$\times$4 × frame count (compared to current SOTA), rich contextual control, and geometric controls. Extensive experiments demonstrate MagicDriveDiT’s ability, unlocking broader applications in autonomous driving. Webpage: flymin.github.io/magicdrivedit/
@@ 1 Introduction
Figure 1: MagicDriveDiT generates high-resolution and long videos with multi-view and control supports, significantly exceeding the limitation of previous works [35, 15, 11, 12].
@@
In this paper, we present MagicDriveDiT, a novel approach based on the DiT architecture, designed to tackle the challenges of high-resolution and long street-view video synthesis with precise control. Building upon DiT [29, 46] integrated with a 3D VAE for geometric control, we introduce an MVDiT block for multi-view video generation, and a novel spatial-temporal conditional encoding for latents derived from CogVAE [42], enabling precise management of geometric controls. Additionally, we enhance contextual description supports by efficiently generating captions for each video. To train the model efficiently, we employ a progressive training strategy, transitioning from short to long videos, which helps the model capture intricate details and generalize to complex scenarios. Furthermore, we utilize videos of diverse resolutions and durations to enhance the model’s generalization capability, enabling it to synthesize longer videos than those used in training.
Our MagicDriveDiT excels in generating highly realistic videos that align with text, road maps, 3D bounding boxes, and different camera perspectives, achieving significantly higher resolution and more frames than previous works (detailed in Table 1). Experiments and comparisons demonstrate the effectiveness of our training and control methods, significantly improving controllable street-view video synthesis. MagicDriveDiT’s flexibility in handling various resolutions, frames, and control signals enables the creation of novel street views suitable for simulations, expanding its potential applications across diverse domains.
@@
  • We design an efficient framework, MagicDriveDiT, leveraging progressive training to achieve high-quality, high-resolution, and long video generation (up to $848\times 1600\times 6$, 241 frames), significantly superior to prior works.
@@
Figure 3: Architecture Overview of MagicDriveDiT. To incorporate different conditions for video generation, MagicDriveDiT adopts a two-branch architecture as in [43] with basic STDiT3 blocks from [6, 46]. We propose the MVDiT block for multi-view consistency and Spatial-Temporal (Box/Traj.) Encoder to inject condition into the Spatial-Temporal (ST) Latent.
@@ 2 Related Work
Video Generation in Autonomous Driving. Video generation is crucial for autonomous driving, with applications in training perception models [11], testing [41], and scene reconstruction [10, 44]. It demands extensive fields of view and dynamic object motion handling, requiring precise controllability [11] and high-resolution video production [21] with more frames [10] and multiple camera perspectives [44]. Diffusion models have improved controllable multi-view video generation, yet existing models [38, 39, 45, 26, 18] lack adequate resolution and frame count for policy testing [16] and data engine applications [11, 10, 44]. By achieving unprecedented resolution and frame count, as compared in Figure 1 and Table 1, MagicDriveDiT addresses the limitations of previous driving video generation methods, enabling more potential applications.
@@
MagicDriveDiT 848$\times$1600$\times$6 241
@@
Diffusion Models and DiT Architectures. Diffusion models [13, 31, 47] generate data by learning the denoising steps from Gaussian noises to samples, widely used in image [8, 30, 37] and video [14] generation. From the modeling’s perspective, Flow Matching [22, 9] simplifies diffusion models, enhancing the efficiency of training and inference. From the implementation’s perspective, the architecture shifts from UNet [13] to DiT, due to better scalability [29, 6], especially for high-resolution tasks [5]. Our MagicDriveDiT also leverages Flow Matching and DiT for scaling to high-resolution and long video generation.
Conditional Generation. Conditional generation is crucial for various applications utilizing generative models. Cross-attention layers from LDM [30] and ControlNet’s [43] additive encoding for grid-shaped control signals are leading methods in controllable diffusion-based generation. In the street-view generation, MagicDrive [11] and MagicDrive3D [10] integrate 3D bounding boxes, BEV maps, ego trajectories, and camera poses for multi-view street scene synthesis. However, these methods are limited to spatial encoding and are not directly applicable to spatial-temporal latents from 3D VAE [42], as shown in Figure 2. MagicDriveDiT presents a novel control paradigm for spatial-temporally compressed latents from 3D VAE, enabling the controllable generation of high-resolution long videos.
@@ 3 Preliminary
LDMs and Flow Matching. For high-resolution image generation, Rombach et al. [30] propose Latent Diffusion Models (LDMs), using a pre-trained VAE for image downsampling and diffusion models for latent generation. This manner is widely adopted in both image generation [6, 9] and video generation [2, 46, 3]. MagicDriveDiT is also based on the VAE+diffusion formulation. With the latest advancement on diffusion models, Esser et al. [9] propose to train large scale diffusion models through simulation-free rectified flow [24, 23] and v-prediction loss [9]:
@@ 4.1 Overview of MagicDriveDiT
Besides, MagicDriveDiT enriches text control beyond previous works (e.g[11]) by generating contextual description through image captioning (Section 4.4). Furthermore, MagicDriveDiT employs progressive training from images to high-resolution short videos, then to long videos, enhancing diffusion model convergence. Using variable-length, multi-resolution videos in training also enables image and video generation at diverse resolutions and frame extrapolation beyond training settings (Section 4.5).
@@ 4.2 Architecture for High-Resolution Long Video
As depicted in Figure 3, MagicDriveDiT utilizes STDiT-3 blocks as per Zheng et al. [46]. The architecture design incorporates two significant modifications. First, to facilitate multi-view generation, the Multi-View DiT (MVDiT) block integrates a cross-view attention layer [11], as shown on the left of Figure 3. Second, given the need for handling multiple control elements, MagicDriveDiT employs cross-attention [30] for text, boxes, camera views, and trajectories, alongside an addictive branch [43] for maps to inject control signals.
@@ 4.3 Spatial-Temporal Control on 3D VAE
Figure 5: MagicDriveDiT supports more diverse text control by enriching the textual description of driving datasets (e.g., nuScene [4], Waymo [33]) than previous works (e.g., [11]). Please see the full videos on our project page.
@@ 4.4 Enrich Text Control by Image Caption
Although geometric descriptors offer precise control over objects and maps, textual descriptions in autonomous driving datasets (e.g., nuScenes) are often restricted to basic weather (sunny/rain) and time (day/night), lacking richer scene context such as road types (e.g., highway/urban) or background elements (e.g., buildings, trees). To enhance textual control, we regenerated video captions using a Multimodal Large Language Model (MLLM) [17]. By prompting MLLM to focus on aspects beyond object categories, trajectories, and road structures, we ensure efficiency and consistency. Specifically, first, it eliminates the need for the MLLM to describe dynamics, thus only the middle frame of each video is used; second, it avoids potential conflicts between multiple control signals. Figure 5 shows the diverse contextual description supports of MagicDriveDiT.
@@ 5.1 Experimental Setups
Dataset and Baselines. We employ the nuScenes dataset [4], a prominent dataset for street-view generation [11, 40, 26], to evaluate MagicDriveDiT. We adhere to the official splits, using 700 multi-view videos for training and 150 for validation. Our primary baseline is MagicDrive [11], comprising three models: MagicDrive (16f) for 16-frame video generation; MagicDrive (60f), which extends MagicDrive (16f) to 60 frames, both from Gao et al. [11]; and MagicDrive3D, a 16-frame model from Gao et al. [10]. Moreover, we employed the Waymo Open Dataset [33] to further fine-tune the model, aiming to validate its scalability and to facilitate the generation of videos with diverse environmental styles and different numbers of perspectives. Further details are provided in Appendix D.
@@
Figure 6: Qualitative Comparison between MagicDriveDiT and MagicDrive [11]. Frames are extracted from the generated videos. To conserve space, we only present the 3 (out of 6) perspectives that include the front view. Two crops from the generated views are magnified and shown on the right. By generating 4$\times$ resolution of MagicDrive, the synthesized street view of our model contains finer details.
@@
Model Setup. In accordance with Section 4.2, we adopt the 3D VAE framework from CogVideoX [42] and train diffusion models for street-view video generation from scratch [^3] For more discussion on this choice, see Section 5.3 and Appendix H.. Initially, the model comprises only spatial blocks for base and control blocks, focusing on image generation. Temporal blocks are incorporated in the second stage, forming the full architecture of MagicDriveDiT. Further training details can be found in Appendix E.
@@ 5.2 Results and Analysis
Generation Quality. MagicDriveDiT excels in both video and image generation tasks. In video tasks, it significantly reduces FVD compared to MagicDrive (Table 2), due to the DiT architecture enhancing inter-frame consistency and spatial-temporal condition encoding for precise control of object motion and positioning. Further illustrated in Figure 6, MagicDriveDiT generates high-resolution videos that not only improve quality but also incorporate more intricate details, closely resembling footage captured by real cameras. This enhancement is achieved through our advanced training on variable lengths and resolutions, allowing for more realistic and detailed outputs. We further conduct human evaluation on consistency in Appendix F.
Beneficial from the mixing training approach, MagicDriveDiT is capable of image generation. As shown in Table 3, MagicDriveDiT matches baseline performance in multi-view street-view tasks and surpasses baselines in vehicle segmentation mIoU and object detection mAP. This demonstrates the strong generalization capabilities of our spatial-temporal condition encoding.
Controllability. The quantitative results presented in Table 2 and 3 demonstrate that the images and videos generated by MagicDriveDiT effectively reflect the control conditions. Additionally, Figure 5 and 8 provide visual results, showing that multiple control conditions can independently influence the generated content. For instance, weather can be altered via text input (from sunny to rainy) while maintaining road structure, and the trajectories of other vehicles and the ego vehicle. By varying the combinations of conditions, MagicDriveDiT is capable of producing a diverse range of high-quality street-view videos.
@@
Figure 8: MagicDriveDiT generates high-resolution (e.g., 424$\times$800 here) street-view videos for 241 frames (i.e., the full length of nuScenes videos, approximately 20 seconds at 12 FPS) with multiple controls (i.e., road map, object boxes, ego trajectory, and text). Notably, the 241-frame length at 424$\times$800 is unseen during training, demonstrating our method’s generalization capability to video length. We annotate the ego-vehicle trajectory and selected objects to aid localization, with same-color boxes denoting the same object. Due to space constraints, the “rainy” example includes only two frames; additional examples can be found in Appendix J and on our project page.
@@ 5.3 Ablation Studies
Figure 9: Validation Loss through Training with Different ST Encodings. 4×4$\times$4 ×down (our methods in MagicDriveDiT) can help the model converge, performing the best among all the encodings.
@@
Variable Length and Resolution Training. As presented in Section 4.5, MagicDriveDiT adopts a training strategy by mixing different length and resolution videos. Our ablation study demonstrates the effectiveness of this strategy. As shown in Table 4, the limitations of VAE (as discussed in Section 5.2) are evident when using only low-resolution videos (17$\times$224$\times$400), indicated by worse FVD and lower mAP and mIoU scores compared to other settings. Incorporating longer videos improves the model’s controllability (both mAP and mIoU are higher), while incorporating high-resolution videos enhances overall quality (all three metrics are notably improved).
@@ 6 Applications
Extrapolation for Longer Video Generation. Through variable length and resolution training (Section 4.5), MagicDriveDiT effectively generates videos exceeding the training setup’s length. Though trained on videos up to 33$\times$848$\times$1600 and 241$\times$224$\times$400, MagicDriveDiT successfully produces 241$\times$848$\times$1600 videos (full length and full resolution of nuScenes), which are 8×8$\times$8 × the length of the training samples, as demonstrated in Figure 8 and videos on our project page. To further validate this capability, we compare the first-16-frame FVD for short videos (17 frames) and per-16-frame FVD for longer videos (65 frames and above). As shown in Table 5, the 16-frame FVDs remain consistent across both seen and extrapolated configurations, extending to 129$\times$848$\times$1600$\times$6 generation (this is not the upper limit; please see note in the Appendix). This affirms the model’s robust generalization capability.
@@
Fast Generalization on Other Datasets. To evaluate the generalization of MagicDriveDiT, we fine-tuned the Stage 3 model on the Waymo dataset [33], enabling rapid generation of 3-view videos in $1$ day ($1$k+ steps) with strong controllability (examples shown in Figure 5 and more videos can be found on our project page). Moreover, mixed training on Waymo and nuScenes datasets further enhanced the model’s ability to generate videos with varying perspectives, improving overall quality. The final model achieved an FVD of 105.17 on Waymo and 74.30 on nuScenes, surpassing results in Table 2.

### v3

# MagicDrive-V2: High-Resolution Long Video Generation for Autonomous Driving with Adaptive Control

@@ Abstract
The rapid advancement of diffusion models has greatly improved video synthesis, especially in controllable video generation, which is vital for applications like autonomous driving. Although DiT with 3D VAE has become a standard framework for video generation, it introduces challenges in controllable driving video generation, especially for geometry control, rendering existing control methods ineffective. To address these issues, we propose MagicDrive-V2, a novel approach that integrates the MVDiT block and spatial-temporal conditional encoding to enable multi-view video generation and precise geometric control. Additionally, we introduce an efficient method for obtaining contextual descriptions for videos to support diverse textual control, along with a progressive training strategy using mixed video data to enhance training efficiency and generalizability. Consequently, MagicDrive-V2 enables multi-view driving video synthesis with $\times3$ resolution and 4×4$\times$4 × frame count (compared to current SOTA), rich contextual control, and geometric controls. Extensive experiments demonstrate MagicDrive-V2’s ability, unlocking broader applications in autonomous driving. Webpage: flymin.github.io/magicdrive-v2/
@@ 1 Introduction
Figure 1: MagicDrive-V2 generates high-resolution and long videos with multi-view and control supports, significantly exceeding the limitation of previous works [35, 15, 11, 12].
@@
In this paper, we present MagicDrive-V2, a novel approach based on the DiT architecture, designed to tackle the challenges of high-resolution and long street-view video synthesis with precise control. Building upon DiT [29, 46] integrated with a 3D VAE for geometric control, we introduce an MVDiT block for multi-view video generation, and a novel spatial-temporal conditional encoding for latents derived from CogVAE [42], enabling precise management of geometric controls. Additionally, we enhance contextual description supports by efficiently generating captions for each video. To train the model efficiently, we employ a progressive training strategy, transitioning from short to long videos, which helps the model capture intricate details and generalize to complex scenarios. Furthermore, we utilize videos of diverse resolutions and durations to enhance the model’s generalization capability, enabling it to synthesize longer videos than those used in training.
Our MagicDrive-V2 excels in generating highly realistic videos that align with text, road maps, 3D bounding boxes, and different camera perspectives, achieving significantly higher resolution and more frames than previous works (detailed in Table 1). Experiments and comparisons demonstrate the effectiveness of our training and control methods, significantly improving controllable street-view video synthesis. MagicDrive-V2’s flexibility in handling various resolutions, frames, and control signals enables the creation of novel street views suitable for simulations, expanding its potential applications across diverse domains.
@@
  • We design an efficient framework, MagicDrive-V2, leveraging progressive training to achieve high-quality, high-resolution, and long video generation (up to $848\times 1600\times 6$, 241 frames), significantly superior to prior works.
@@
Figure 3: Architecture Overview of MagicDrive-V2. To incorporate different conditions for video generation, MagicDrive-V2 adopts a two-branch architecture as in [43] with basic STDiT3 blocks from [6, 46]. We propose the MVDiT block for multi-view consistency and Spatial-Temporal (Box/Traj.) Encoder to inject condition into the Spatial-Temporal (ST) Latent.
@@ 2 Related Work
Video Generation in Autonomous Driving. Video generation is crucial for autonomous driving, with applications in training perception models [11], testing [41], and scene reconstruction [10, 44]. It demands extensive fields of view and dynamic object motion handling, requiring precise controllability [11] and high-resolution video production [21] with more frames [10] and multiple camera perspectives [44]. Diffusion models have improved controllable multi-view video generation, yet existing models [38, 39, 45, 26, 18] lack adequate resolution and frame count for policy testing [16] and data engine applications [11, 10, 44]. By achieving unprecedented resolution and frame count, as compared in Figure 1 and Table 1, MagicDrive-V2 addresses the limitations of previous driving video generation methods, enabling more potential applications.
@@
MagicDrive-V2 848$\times$1600$\times$6 241
@@
Diffusion Models and DiT Architectures. Diffusion models [13, 31, 47] generate data by learning the denoising steps from Gaussian noises to samples, widely used in image [8, 30, 37] and video [14] generation. From the modeling’s perspective, Flow Matching [22, 9] simplifies diffusion models, enhancing the efficiency of training and inference. From the implementation’s perspective, the architecture shifts from UNet [13] to DiT, due to better scalability [29, 6], especially for high-resolution tasks [5]. Our MagicDrive-V2 also leverages Flow Matching and DiT for scaling to high-resolution and long video generation.
Conditional Generation. Conditional generation is crucial for various applications utilizing generative models. Cross-attention layers from LDM [30] and ControlNet’s [43] additive encoding for grid-shaped control signals are leading methods in controllable diffusion-based generation. In the street-view generation, MagicDrive [11] and MagicDrive3D [10] integrate 3D bounding boxes, BEV maps, ego trajectories, and camera poses for multi-view street scene synthesis. However, these methods are limited to spatial encoding and are not directly applicable to spatial-temporal latents from 3D VAE [42], as shown in Figure 2. MagicDrive-V2 presents a novel control paradigm for spatial-temporally compressed latents from 3D VAE, enabling the controllable generation of high-resolution long videos.
@@ 3 Preliminary
LDMs and Flow Matching. For high-resolution image generation, Rombach et al. [30] propose Latent Diffusion Models (LDMs), using a pre-trained VAE for image downsampling and diffusion models for latent generation. This manner is widely adopted in both image generation [6, 9] and video generation [2, 46, 3]. MagicDrive-V2 is also based on the VAE+diffusion formulation. With the latest advancement on diffusion models, Esser et al. [9] propose to train large scale diffusion models through simulation-free rectified flow [24, 23] and v-prediction loss [9]:
@@ 4.1 Overview of MagicDrive-V2
Besides, MagicDrive-V2 enriches text control beyond previous works (e.g[11]) by generating contextual description through image captioning (Section 4.4). Furthermore, MagicDrive-V2 employs progressive training from images to high-resolution short videos, then to long videos, enhancing diffusion model convergence. Using variable-length, multi-resolution videos in training also enables image and video generation at diverse resolutions and frame extrapolation beyond training settings (Section 4.5).
@@ 4.2 Architecture for High-Resolution Long Video
As depicted in Figure 3, MagicDrive-V2 utilizes STDiT-3 blocks as per Zheng et al. [46]. The architecture design incorporates two significant modifications. First, to facilitate multi-view generation, the Multi-View DiT (MVDiT) block integrates a cross-view attention layer [11], as shown on the left of Figure 3. Second, given the need for handling multiple control elements, MagicDrive-V2 employs cross-attention [30] for text, boxes, camera views, and trajectories, alongside an addictive branch [43] for maps to inject control signals.
@@ 4.3 Spatial-Temporal Control on 3D VAE
Figure 5: MagicDrive-V2 supports more diverse text control by enriching the textual description of driving datasets (e.g., nuScene [4], Waymo [33]) than previous works (e.g., [11]). Please see the full videos on our project page.
@@ 4.4 Enrich Text Control by Image Caption
Although geometric descriptors offer precise control over objects and maps, textual descriptions in autonomous driving datasets (e.g., nuScenes) are often restricted to basic weather (sunny/rain) and time (day/night), lacking richer scene context such as road types (e.g., highway/urban) or background elements (e.g., buildings, trees). To enhance textual control, we regenerated video captions using a Multimodal Large Language Model (MLLM) [17]. By prompting MLLM to focus on aspects beyond object categories, trajectories, and road structures, we ensure efficiency and consistency. Specifically, first, it eliminates the need for the MLLM to describe dynamics, thus only the middle frame of each video is used; second, it avoids potential conflicts between multiple control signals. Figure 5 shows the diverse contextual description supports of MagicDrive-V2.
@@ 5.1 Experimental Setups
Dataset and Baselines. We employ the nuScenes dataset [4], a prominent dataset for street-view generation [11, 40, 26], to evaluate MagicDrive-V2. We adhere to the official splits, using 700 multi-view videos for training and 150 for validation. Our primary baseline is MagicDrive [11], comprising three models: MagicDrive (16f) for 16-frame video generation; MagicDrive (60f), which extends MagicDrive (16f) to 60 frames, both from Gao et al. [11]; and MagicDrive3D, a 16-frame model from Gao et al. [10]. Moreover, we employed the Waymo Open Dataset [33] to further fine-tune the model, aiming to validate its scalability and to facilitate the generation of videos with diverse environmental styles and different numbers of perspectives. Further details are provided in Appendix D.
@@
Figure 6: Qualitative Comparison between MagicDrive-V2 and MagicDrive [11]. Frames are extracted from the generated videos. To conserve space, we only present the 3 (out of 6) perspectives that include the front view. Two crops from the generated views are magnified and shown on the right. By generating 4$\times$ resolution of MagicDrive, the synthesized street view of our model contains finer details.
@@
Model Setup. In accordance with Section 4.2, we adopt the 3D VAE framework from CogVideoX [42] and train diffusion models for street-view video generation from scratch [^3] For more discussion on this choice, see Section 5.3 and Appendix H.. Initially, the model comprises only spatial blocks for base and control blocks, focusing on image generation. Temporal blocks are incorporated in the second stage, forming the full architecture of MagicDrive-V2. Further training details can be found in Appendix E.
@@ 5.2 Results and Analysis
Generation Quality. MagicDrive-V2 excels in both video and image generation tasks. In video tasks, it significantly reduces FVD compared to MagicDrive (Table 2), due to the DiT architecture enhancing inter-frame consistency and spatial-temporal condition encoding for precise control of object motion and positioning. Further illustrated in Figure 6, MagicDrive-V2 generates high-resolution videos that not only improve quality but also incorporate more intricate details, closely resembling footage captured by real cameras. This enhancement is achieved through our advanced training on variable lengths and resolutions, allowing for more realistic and detailed outputs. We further conduct human evaluation on consistency in Appendix F.
Beneficial from the mixing training approach, MagicDrive-V2 is capable of image generation. As shown in Table 3, MagicDrive-V2 matches baseline performance in multi-view street-view tasks and surpasses baselines in vehicle segmentation mIoU and object detection mAP. This demonstrates the strong generalization capabilities of our spatial-temporal condition encoding.
Controllability. The quantitative results presented in Table 2 and 3 demonstrate that the images and videos generated by MagicDrive-V2 effectively reflect the control conditions. Additionally, Figure 5 and 8 provide visual results, showing that multiple control conditions can independently influence the generated content. For instance, weather can be altered via text input (from sunny to rainy) while maintaining road structure, and the trajectories of other vehicles and the ego vehicle. By varying the combinations of conditions, MagicDrive-V2 is capable of producing a diverse range of high-quality street-view videos.
@@
Figure 8: MagicDrive-V2 generates high-resolution (e.g., 424$\times$800 here) street-view videos for 241 frames (i.e., the full length of nuScenes videos, approximately 20 seconds at 12 FPS) with multiple controls (i.e., road map, object boxes, ego trajectory, and text). Notably, the 241-frame length at 424$\times$800 is unseen during training, demonstrating our method’s generalization capability to video length. We annotate the ego-vehicle trajectory and selected objects to aid localization, with same-color boxes denoting the same object. Due to space constraints, the “rainy” example includes only two frames; additional examples can be found in Appendix J and on our project page.
@@ 5.3 Ablation Studies
Figure 9: Validation Loss through Training with Different ST Encodings. 4×4$\times$4 ×down (our methods in MagicDrive-V2) can help the model converge, performing the best among all the encodings.
@@
Variable Length and Resolution Training. As presented in Section 4.5, MagicDrive-V2 adopts a training strategy by mixing different length and resolution videos. Our ablation study demonstrates the effectiveness of this strategy. As shown in Table 4, the limitations of VAE (as discussed in Section 5.2) are evident when using only low-resolution videos (17$\times$224$\times$400), indicated by worse FVD and lower mAP and mIoU scores compared to other settings. Incorporating longer videos improves the model’s controllability (both mAP and mIoU are higher), while incorporating high-resolution videos enhances overall quality (all three metrics are notably improved).
@@ 6 Applications
Extrapolation for Longer Video Generation. Through variable length and resolution training (Section 4.5), MagicDrive-V2 effectively generates videos exceeding the training setup’s length. Though trained on videos up to 33$\times$848$\times$1600 and 241$\times$224$\times$400, MagicDrive-V2 successfully produces 241$\times$848$\times$1600 videos (full length and full resolution of nuScenes), which are 8×8$\times$8 × the length of the training samples, as demonstrated in Figure 8 and videos on our project page. To further validate this capability, we compare the first-16-frame FVD for short videos (17 frames) and per-16-frame FVD for longer videos (65 frames and above). As shown in Table 5, the 16-frame FVDs remain consistent across both seen and extrapolated configurations, extending to 129$\times$848$\times$1600$\times$6 generation (this is not the upper limit; please see note in the Appendix). This affirms the model’s robust generalization capability.
@@
Fast Generalization on Other Datasets. To evaluate the generalization of MagicDrive-V2, we fine-tuned the Stage 3 model on the Waymo dataset [33], enabling rapid generation of 3-view videos in $1$ day ($1$k+ steps) with strong controllability (examples shown in Figure 5 and more videos can be found on our project page). Moreover, mixed training on Waymo and nuScenes datasets further enhanced the model’s ability to generate videos with varying perspectives, improving overall quality. The final model achieved an FVD of 105.17 on Waymo and 74.30 on nuScenes, surpassing results in Table 2.